# FM2DS: Few-Shot Multimodal Multihop Data Synthesis with Knowledge Distillation for Question Answering

### v2

@@ Abstract
Multimodal multihop question answering is a complex task that requires reasoning over multiple sources of information, such as images and text, to answer questions. While there has been significant progress in visual question answering, the multihop setting remains unexplored due to the lack of high-quality datasets. Current methods focus on single-hop question answering or a single modality, which makes them unsuitable for real-world scenarios such as analyzing multimodal educational materials, summarizing lengthy academic articles, or interpreting scientific studies that combine charts, images, and text. To address this gap, we propose a novel methodology, introducing the first framework for creating a high-quality dataset that enables training models for multimodal multihop question answering. Our approach consists of a 5-stage pipeline that involves acquiring relevant multimodal documents from Wikipedia, synthetically generating high-level questions and answers, and validating them through rigorous criteria to ensure quality data. We evaluate our methodology by training models on our synthesized dataset and testing on two benchmarks, our results demonstrate that, with an equal sample size, models trained on our synthesized data outperform those trained on human-collected data by 1.9 in exact match (EM) on average. We believe our data synthesis method will serve as a strong foundation for training and evaluating multimodal multihop question answering models [^1] Code and data are available at https://github.com/ServiceNow/FM2DS.
@@ 1 Introduction
Figure 1: In contrast to traditional datasets that depend on human annotators, templates, and information snippets as sources, FM2DS is a fully automated approach that utilizes complete documents as its sources. FM2DS incorporates validation steps to ensure that the generated questions are answerable, multimodal, and multihop.
Few-shot learning for multimodal multihop question answering (MMQA) aims to address complex queries by reasoning across multiple modalities, including text and images. This approach is crucial in scenarios where questions require reasoning and integrating information across long texts, images, tables, and other modalities. Recent advancements in this area highlight the potential of large language models (LLMs) and large vision language models (LVLMs) to enhance reasoning capabilities [38, 20, 6], yet existing frameworks often rely heavily on extensive human-annotated datasets, which is not always feasible for practical applications[35].
@@
Current methodologies in MMQA typically leverage in-context learning methods, prompting LVLMs to retrieve relevant information from multimodal sources [36] and then perform reasoning [41]. However, these models often demand significant computational resources due to their large parameter counts, making them costly to deploy even during inference [42]. This limitation emphasizes the need for more efficient frameworks that can operate effectively with minimal annotated data. A practical solution is to use a smaller model capable of retrieving the necessary information from sources before performing reasoning. This can be achieved by fine-tuning the model on a MMQA dataset. However, existing datasets often consist of short information snippets rather than full documents [5], or they rely on repetitive question templates [35], which limits the model’s generalizability. They also primarily address unimodal or simpler multimodal tasks that do not require long texts with multiple images [21, 7]. Additionally, creating new similar datasets is challenging, requiring extensive human annotation [30, 10].
@@
In this work, we propose FM2DS, a novel data synthesis framework designed specifically for few-shot multimodal multihop question answering. Our approach synthesizes MMQA data from documents that are interconnected through various relationships, such as thematic similarities or sequential events. This framework leverages naturally occurring document relationships and requires minimal hand-crafted data, thereby broadening the range of reasoning types employed in question generation. As illustrated in Figure 1, FM2DS enables the generation of non-templated question-answer pairs based on full documents rather than brief information snippets. By integrating a query component that offers a step-by-step guide for retrieving relevant information from various documents, this framework enables smaller models to handle complex questions similarly to larger models. This methodology allows users to create a custom MMQA dataset with fewer than ten human-annotated samples, thereby facilitating the fine-tuning of a smaller vision-language model (VLM) for specific applications.
In our method, we utilize Wikipedia as a primary data source due to its extensive knowledge base and hyperlink structure. Next, by using document pairs with shared topical relevance or hyperlink connections, we prompt LVLMs to perform multiple generation tasks, including question generation, question answering, and query generation, similar to previous work in text-only settings [13, 8]. By leveraging few-shot samples and tapping into the knowledge and reasoning capabilities of LVLMs, we generate diverse question-answer pairs, accompanied by queries, to streamline and accelerate the learning process for smaller VLMs.
Through empirical evaluation on established benchmarks for MMQA, we show that FM2DS significantly enhances model performance across various configurations. By fine-tuning smaller VLMs on our synthesized data, we achieve superior performance on human-annotated test sets compared to fine-tuning the same model on the human-annotated data. This work advances the state-of-the-art in MMQA by introducing a scalable framework that harnesses the potential of few-shot learning. By addressing the critical gap between single-hop and multi-hop reasoning in multimodal contexts, our work presents a novel approach that significantly reduces the dependency on extensively annotated datasets while maintaining robust performance. The proposed methodology not only bridges the existing limitations in MMQA systems but also establishes a foundation for more efficient and adaptable question answering systems that can reason across diverse modalities and multiple inference steps. The key contributions of this work are:
  • Introducing a novel framework for synthesizing high-quality training data for VLMs to effectively perform Multimodal Multi-hop Question Answering (MMQA).
  • Creating a robust validation pipeline that verifies the quality of generated questions, answers and the corresponding queries, thus maintaining high quality of the synthesized data.
  • Presenting a challenging MMQA benchmark that can be utilized to test the performance of models on questions that require reasoning over multiple modalities and information sources to answer, reflecting the real-world setting.
  • Through our empirical evaluation, we demonstrate that models trained on our synthetically generated dataset achieve higher performance over the models trained on manually curated datasets.
@@ 2 Related Work
Within the Question Answering (QA) literature, synthesis of training data has been predominantly focused on unimodal (text-only) scenarios. We review various seminal works that have established the foundation for our work in few-shot data synthesis.
Figure 2: The Five-Stage Pipeline for FM2DS. First we retrieve relevant documents from Wikipedia to create a pool of related documents based on hyperlinks and topics (Stage 1). In Stage2, we select the few-shot samples from MultiModalQA (MMQA in the figure) [35]. Stage 3 focuses on generating and validating questions to make sure they are answerable, multihop, and multimodal . In Stage 4, answers are generated and validated. Finally, in Stage 5 we generate queries related to the documents, which are also validated to ensure relevance and accuracy.
@@
Synthetic data has become a widely used resource for model training. He et al. [16] propose a general framework that utilizes synthetic text generated by language models (LMs) for knowledge distillation, self-training, and few-shot learning, demonstrating that combining labeled and pseudo-labeled synthetic data can lead to notable performance gains in NLP tasks. In other cases, rather than synthetic text alone, entire datasets are generated for specific tasks such as classification [37]. For instance, Li et al. [26] evaluate the effectiveness of LLMs, specifically GPT-3.5, in producing synthetic datasets for text classification, offering insights into the consistency and reliability of LLM-generated data. In another work, Chen et al. [8] present a framework that significantly improves the performance of smaller language models on multi-hop QA tasks, even with minimal human-annotated data.
@@
Research on multimodal data synthesis using LVLMs is still underexplored, most work focuses on creating entirely new data based on knowledge encoded in these models. For instance, Zhang et al. [43] design a multimodal benchmark utilizing large vision and language models and their code capabilities to synthesize massive abstract images and visual reasoning instructions across daily scenarios. In another study, Mehta et al. [32] tackles the scarcity of training data for models that generate both speech and gestures, using unimodal synthesis models to create synthetic multimodal training data, which is then employed to pre-train a joint synthesis model. In the field of multimodal question answering, Wu et al. [39] introduced SMMQG, a method that generates questions for multimodal documents using multimodal RAG. In this work, we present a novel approach for synthetically generating a question answering dataset, leveraging the reasoning capabilities of LVLMs to create not only multimodal questions but also multihop ones, bringing them closer to real-world scenarios.
@@ 3 Proposed Method: FM2DS
Our five-stage pipeline for FM2DS shown in Figure 2 is designed to synthesize high-quality multimodal question-answer pairs. The process begins with retrieving relevant documents through topic matching and Wikipedia hyperlinks. The subsequent stages involve few-shot sample selection, synthesizing and validating multihop questions, generating corresponding answers, and constructing relevant document queries, with each stage incorporating specific validation mechanisms. We elaborate on individual components in the following sections. Refer to Appendix E for examples of generated data by FM2DS.
@@ 3.1 Stage 1: Creating a Pool of Related Documents
We collect a pool of relevant documents using Wikipedia pages leveraging Wikipedia’s vast repository of interconnected information. For this, we used the WikiWeb2M dataset [4], which contains nearly 2 million Wikipedia pages, providing a comprehensive source of information. These documents are linked through two complementary methods: existing hyperlinks and topics extracted using multimodal topic modeling with Multimodal-Contrast model [15]. The hyperlink structure provides explicit, human-curated connections between related concepts, while the multimodal topic modeling technique extracts latent thematic relationships from both textual and non-textual content. This dual approach ensures a thorough compilation of related information, capturing both obvious and nuanced connections across documents, while integrating both visual and textual modalities.
@@ 3.2 Stage 2: Creating Few-Shot Samples
To generate questions, we leverage existing multimodal multihop question-answering dataset MultiModalQA [35] that require reasoning across various modalities, including text, images, and tables. For our purposes, we utilized a custom subset of the MultiModalQA train set. Using the links to Wiikpedia pages provided in the dataset, we create a pool for few-shot samples that include the full Wikipedia pages —complete with images and tables in HTML format—along with the question. In our experiments, we randomly select up to three of these samples for question generation.
@@ 3.3 Stage 3: Question Generation and Validation
We use GPT-4-turbo [33] to generate questions based on few-shot samples created from the MultiModalQA dataset. Due to the model’s context length limitations, we restricted the input to two or three related documents. Our prompt, detailed in Appendix A, query the model to produce multimodal, multihop questions that required input from at least two modalities and drew on information from all provided documents. For instance, if two documents were given, the question needed both to be answerable; if three documents are provided, it can utilize any two or all three of them. We explicitly instructed the model that simply concatenating two independent questions with an “and” to reference multiple documents would not qualify as a valid multihop question. For instance How did Albert Einstein contribute to the Theory of Relativity and when Princeton established? is not a true multihop question, even though it involves two documents (Einstein’s death in Princeton being the link). A genuine multihop question requires integrating information across sources to necessitate reasoning over multiple documents.
@@
Our framework incorporates multiple validation layers to verify that synthesized questions satisfy the requisite criteria for both multihop reasoning complexity and multimodal integration. Although the model was instructed during question generation to avoid creating questions that simply concatenate independent questions answerable by a single document, we evaluated this further. For this, we employed LLama-3.1-8B [2] to break each question down into simpler components. We then assessed whether each component could be answered using only one document. If all parts of the question were solvable with a single document, we discarded such question that include unrelated facts. Here is an example of question with unrelated facts:
Example (Unrelated Facts): In what year did Mike Tyson become the youngest heavyweight champion, and who is the president of the United States?
Alternatively, we kept only the initial component that required information from multiple documents to ensure the new question met the multihop criteria. One potential issue with this approach was that, although the facts in the questions are related, they often led to open-ended questions where the answer needed to be an explanation or a combination of two responses. For example:
Example (Related Facts, Open-ended): In what year did Mike Tyson become the youngest heavyweight champion, and who was the president of the United States at that time?
In order to follow the standard of question answering, and make the evaluation process easier, we used GPT-4o to rephrase the question without conjunctions while maintaining its multihop nature, resulting in:
Example(Concise Multihop Query): Who was the president of the United States when Mike Tyson became the youngest heavyweight champion?
Another critical part of validation was ensuring the multimodal nature of the questions. After confirming that a question was multihop, we tested whether it could still be answered if the documents were reduced to a single modality (e.g., text-only, image-only, or table-only). We used GPT-4o to see if the question is answerable with only one modality (refer to Section 3.4 for more details on how to check if a question is answerable). If a question could be answered using just one modality, we discarded it, as it did not meet the multimodal requirement. This additional check allowed us to refine the dataset, ensuring that the final set of questions required reasoning over multiple modalities and documents, as intended.
@@ 3.4 Stage 4: Answer Generation and Validation
For answer generation, we used GPT-4o to produce concise answers based on multiple documents, which included both textual and visual information such as images. The model was explicitly instructed to answer the question in the shortest way possible, providing only the key information without offering any additional explanation. To help the model focus on specific details of images in the given documents to answer the multimodal question, we include question-related captions for the images. For example, if the question asks about the geometric shapes in an image (see Figure 11), the model generates a caption describing the shapes. This makes it easier for the model to answer the question accurately.
@@
We validated the generated answers using information extraction techniques, specifically named entity recognition (NER) and relation extraction, as they have been previously used for validating answers [34, 14]. NER identified key named entities (e.g., persons, organizations, locations) and numerical data in both the answer and the documents, which were compared to ensure alignment with the source content. Relation extraction was used to verify whether the relationships between entities in the generated answer matched those in the documents. By ensuring the correctness of both named entities and their relations, we validated that the answers were accurate, concise, and consistent with the multimodal data provided  [^2] We utilized Spacy [40] for relation extraction.. For including the images in the process, we used captions generated by GPT-4o. Specifically, based on what the question requires from the image (e.g., the color of a building), the caption is generated to include that information. Additionally, to mitigate hallucinations, we prompted GPT-4o five times, and if the model generated the same correct response across all five prompts, we accepted it. To evaluate the effectiveness of our answer validation step, we designed a human evaluation where the result can be found in Section 6.
@@ 3.5 Stage 5: Query Generation and Validation
We create queries based on the question-answer pairs and the related documents to facilitate a more effective retrieval process. These queries are meticulously designed to guide the search for relevant information, allowing us to pinpoint specific data points that are crucial for formulating accurate answers. This strategic approach helped us narrow down the information within the documents, enabling us to identify and extract key details such as entities, relationships and contextual information that align with the questions. This targeted querying facilitated an effective information extraction approach, that the generated answers are comprehensive and directed supported by the evidence found in the documents.
@@
To validate the queries, we used multimodal retrieval-augmented generator (MuRAG) [9], which integrates both textual and visual information by encoding documents and images into a shared embedding space. This approach allows us to retrieve relevant documents based on both text and visual context, ensuring a more comprehensive validation. The validation process involves querying the MuRAG database and retrieving the top-5 most relevant documents for each generated query. If more than one source document appears in the top-5 results across all queries, it indicates that the query is well-formed and effectively captures diverse, relevant information from multiple sources. Moreover, this validation ensures that the queries are correct and serves as a guide for teaching a smaller model how to retrieve the necessary information to answer the questions.
4 Proposed Benchmark: M2QA-Bench
We also propose a benchmark, M2QA-Bench, to assess the LVLMs performance on a more complicated MMQA task with full documents. M2QA-Bench consists of 500 Q&A pairs, each designed to challenge the model’s ability to perform a complex reasoning task. The questions are not templated into a specific structure (as in some existing works [35]), instead, they are diverse and challenging. Additionally, answering the questions require access to full documents, where both information extraction and reasoning across different modalities (e.g., images and tables) are essential. Table 1 presents the key statistics of M2QA-Bench. Refer to Appendix E for benchmark samples created using FM2DS.
@@ 4 Proposed Benchmark: M2QA-Bench
To create this benchmark, we used the FM2DS pipeline to generate 1,000 samples which were further verified by three human-annotators on whether it was correct, multihop, and multimodal, as well as whether the generated answer was accurate. Samples without issues received a score of 1, while those with issues received a score of 0. After completing the annotation process, we removed samples with an average score below 1 and randomly selected 500 samples from the remaining set.

### v4

# with Knowledge Distillation for Question Answering

@@ Abstract
Multimodal multihop question answering (MMQA) requires reasoning over images and text from multiple sources. Despite advances in visual question answering, this multihop setting remains underexplored due to a lack of quality datasets. Existing methods focus on single-hop, single-modality, or short texts, limiting real-world applications like interpreting educational documents with long, multimodal content. To fill this gap, we introduce FM2DS, the first framework for creating a high-quality dataset for MMQA. Our approach consists of a 5-stage pipeline that involves acquiring relevant multimodal documents from Wikipedia, synthetically generating high-level questions and answers, and validating them through rigorous criteria to ensure data quality. We evaluate our methodology by training models on our synthesized dataset and testing on two benchmarks: MultimodalQA and WebQA. Our results demonstrate that, with an equal sample size, models trained on our synthesized data outperform those trained on human-collected data by 1.9 in exact match (EM) score on average. Additionally, we introduce M2QA-Benchwith 1k samples, the first benchmark for MMQA on long documents, generated using FM2DSand refined by human annotators. We believe our data synthesis method will serve as a strong foundation for training and evaluating MMQA models  [^1] Code and benchmark are available at: https://github.com/ServiceNow/FM2DS.
@@ 1 Introduction
Figure 1: Unlike traditional datasets that rely on human annotators, templates, or snippets, FM2DS is fully automated, using long documents as sources and applying validation to ensure questions are answerable, multimodal, and multihop.
@@
Multimodal multihop question answering (MMQA) involves answering complex questions by integrating information from text, images, and tables. In real-world applications such as interpreting medical documents, this challenge is amplified by the need to reason over long, multimodal content. Current methodologies in MMQA typically leverage in-context learning methods, prompting LVLMs to retrieve relevant information from multimodal sources (Tejaswi et al., 2024) and then perform reasoning (Yang et al., 2023). However, these models often demand significant computational resources due to their large parameter counts, making them costly to deploy even during inference (Ye et al., 2024). This limitation emphasizes the need for more efficient frameworks that can operate effectively with minimal annotated data. A practical solution is to use a smaller model capable of both retrieving the necessary information from sources and performing reasoning. This can be achieved by fine-tuning the model on a MMQA dataset. Existing datasets often rely on short snippets or repetitive templates, limiting generalizability to complex settings with long texts and multiple modalities (Chang et al., 2021; Talmor et al., 2021; Jiang et al., 2024; Chen et al., 2024a). Additionally, creating new similar datasets is challenging, requiring extensive human annotation (Lu et al., 2022b; Chen et al., 2023a).
@@
In this work, we propose FM2DS, a novel data synthesis framework designed specifically for MMQA over long documents. Our approach synthesizes MMQA data from documents that are interconnected through various relationships, such as thematic similarities or sequential events. This framework leverages naturally occurring document relationships and requires minimal hand-crafted data, thereby broadening the range of reasoning types used in question generation. As illustrated in Figure 1, FM2DS enables the generation of non-templated question-answer pairs based on full documents rather than brief information snippets. The data generated by our method FM2DS includes query component— a step-by-step guide for retrieving relevant information from multiple documents — enabling smaller models trained on this synthesized data to learn how to tackle complex questions in a manner similar to larger models. This methodology allows users to create a custom MMQA dataset with fewer than ten human-annotated samples, thereby facilitating the fine-tuning of smaller LVLMs for specific applications.
FM2DS leverages Wikipedia’s extensive knowledge base and hyperlink structure, to select document pairs with shared topical relevance or hyperlink connections and prompt LVLMs to perform question generation, question answering, and query generation. We incorporate validation steps to enhance the quality of the generated data and discard any outputs that are factually incorrect. Through empirical evaluation on established MMQA benchmarks, we show that FM2DS significantly improves model performance, achieving on average a 1.9 exact match (EM) score improvement across two benchmarks: MultimodalQA and WebQA. Our key contributions are: (I) introducing a new framework for synthesizing high-quality MMQA training data for VLMs; (II) using a robust validation pipeline to ensure data quality; (III) introducing a challenging MMQA benchmark requiring reasoning over multiple modalities and sources; and (IV) showing that models fine-tuned on our synthetic data outperform those trained on human-labeled datasets, advancing MMQA while reducing manual effort.
@@ 2 Related Work
Within the Question Answering (QA) literature, synthesis of training data has been predominantly focused on unimodal (text-only) scenarios. We review various similar works that have established the foundation for our work in few-shot data synthesis.
@@
Synthetic data is increasingly used for model training. He et al. (2022) show that combining labeled and synthetic text from language models (LMs) improves NLP performance. Entire synthetic datasets have also been created for tasks like classification (Tsui, 2024), with Li et al. (2023) demonstrating GPT-3.5’s effectiveness in generating reliable classification data. Similarly, Chen et al. (2024b) show that synthetic data can significantly boost small models on multi-hop QA with minimal human annotation.
@@
Figure 2: The five-stage pipeline for FM2DS begins with retrieving related Wikipedia documents based on hyperlinks and topics (Stage 1). Next, we select few-shot examples from MultiModalQA Talmor et al. (2021) (Stage 2). Stage 3 generates and validates questions to ensure they are answerable, multihop, and multimodal. In Stage 4, answers are generated and validated. Finally, Stage 5 produces and validates queries to guide complex MMQA.
Research on multimodal data synthesis with LVLMs remains limited, with most efforts focused on generating new data from model’s pre-trained knowledge. Zhang et al. (2024) synthesize abstract images with reasoning tasks, while Mehta et al. (2024) generate multimodal data using unimodal models for pre-training. In MMQA, Wu et al. (2024) propose SMMQG, which uses multimodal RAG to generate questions from short snippets, focusing on multimodality rather than multihop reasoning. In contrast, FM2DS uses full multimodal documents, resulting in a more challenging dataset with diverse multihop questions that better reflect real-world tasks. Moreover, while SMMQG is confined to predefined question types, FM2DS enables large-scale generation and supports knowledge distillation for smaller models through step-by-step queries that guide complex multi-document reasoning.
@@ 3 Proposed Method: FM2DS
Our five-stage pipeline for FM2DS (Figure 2) synthesizes high-quality multimodal QA pairs. It begins by grouping documents via topic matching and Wikipedia hyperlinks, followed by few-shot sample selection, question synthesis, answer generation, and query construction, each with their built-in validation. See Appendix I for examples.
@@ 3.1 Stage 1: Creating a Pool of Related Documents
We collect relevant documents from Wikipedia using the WikiWeb2M dataset (Burns et al., 2023), which includes nearly 2 million pages. Documents are linked via two methods: hyperlinks and latent topics identified through multimodal topic modeling with the Multimodal-Contrast model (González-Pizarro & Carenini, 2024). Since Multimodal-Contrast can’t handle long documents, we split each document into shorter segments containing at most one image, applied topic modeling to each segment, then merged the results and removed duplicates. This combination captures both clear and subtle relationships across documents, integrating textual and visual information.
@@ 3.2 Stage 2: Creating Few-Shot Samples
We sample multihop questions from MultiModalQA dataset (Talmor et al., 2021), which requires reasoning across text, images, and tables. As our samples are based on full documents rather than short information snippets like in MultimodalQA, we crawled the complete Wikipedia HTML pages using the entity links provided in MultimodalQA, which are associated with the dataset’s examples. We then compiled few-shot samples using these full HTML pages—complete with images and tables—paired with their corresponding questions. We randomly select up to three samples for question generation in our experiments.
@@ 3.3 Stage 3: Question Generation and Validation
We use GPT-4-turbo (OpenAI et al., 2024) to generate multihop, multimodal questions from few-shot samples based on MultiModalQA few-shot examples. Due to context limitations, inputs are limited to grouped sets containing 2 or 3 documents. Our prompt (Appendix A) is designed to ensure that questions require reasoning across all documents and at least two modalities, avoiding unrelated combinations such as: “How did Einstein contribute to relativity and when was Princeton established?”—a question that spans multiple documents but lacks meaningful multihop reasoning.
Question Type
In what year did Mike Tyson become the youngest heavyweight champion, and who is the president of the United States? Unrelated Facts
Question Type
In what year did Mike Tyson become the youngest heavyweight champion, and who was the president of the United States at that time? Related Facts, Open-ended
Question Type
Who was the president of the United States when Mike Tyson became the youngest heavyweight champion? Concise Multihop Question
Table 1: Examples of factual questions with varying degrees of relevance and conciseness, demonstrating progression from unrelated to concise multihop reasoning.
@@
Our framework includes validation stages to ensure questions meet multihop and multimodal criteria. While the model was prompted to avoid simple concatenations, we further evaluated this aspect.
We used LLama-3.1-8B (Abhimanyu Dubey et al., 2024) to decompose questions and check if parts could be answered with a single document. If all parts of the question were answerable with a single document, we discarded such question that include unrelated facts (see Unrelated Facts example in Table 1). Alternatively, we retained only the initial component that required information from multiple documents to ensure the revised question met the multihop criteria. However, a potential issue was that, even when the facts were related, the questions could still become open-ended, requiring explanations or combined answers (see example Related Facts, Open-ended in Table 1). In order to follow the standard of question answering, and make the evaluation process easier, we used GPT-4o to rephrase the question without conjunctions while maintaining its multihop nature, resulting in Concise Multihop Question (Table 1).
Another key step in validation was ensuring the questions were truly multimodal. After verifying that a question was multihop, we tested whether it remained answerable when the documents were limited to a single modality (e.g., text-only, image-only, or table-only). Using GPT-4o (refer to Section 3.4 for details), we checked if the question could be answered with just one modality. If so, we discarded it, as it failed to meet the multimodal requirement. This step helped refine the dataset to include only questions that genuinely required reasoning across multiple modalities and documents.
@@ 3.4 Stage 4: Answer Generation and Validation
We used GPT-4o to generate concise answers from multiple documents, including text and images. The model was instructed to provide a long answer and a short answer with only key information and no extra explanation. To help the model focus on specific details of images in the given documents to answer the multimodal question, we include question-related captions for the images. For example, if the question asks about the geometric shapes in an image (see Figure 15), the model generates a caption describing the shapes. This makes it easier for the model to answer the question accurately.
@@
We validated answers using named entity recognition (NER) and relation extraction, following prior work (Rajpurkar et al., 2018; Fabbri et al., 2022). NER ensured key entities and numbers in the answer matched the documents, while relation extraction verified that entity relationships were consistent with the source content (via Spacy (Wu & He, 2019)). For including image content, we used the same question-related caption generated by GPT-4o (e.g., noting a building’s color if relevant to the question) similar to answer generation. To reduce hallucinations, we prompted GPT-4o five times and accepted answers only if all outputs (5/5) agreed. To evaluate the effectiveness of our answer validation process, we conducted a human study to assess the quality of the filtered questions and answers. The results of this evaluation are presented in Section 6.
@@ 3.5 Stage 5: Query Generation and Validation
We generate queries using GPT-4o based on the question-answer pairs and related documents to enhance retrieval effectiveness. These queries help focus the search on specific, relevant information needed to answer the questions accurately. By narrowing down the content, we can extract key details such as named entities, relationships, and contextual cues aligned with the question. This targeted approach ensures that the generated answers are not only concise and accurate but also directly grounded in evidence from the documents.
@@
To validate the queries, we used MuRAG (Chen et al., 2022), which encodes text and images into a shared embedding space for multimodal retrieval. For each generated query, we retrieved the top-5 documents retrieved by MuRAG. If more than one of the original source documents used to generate the question appeared in the top-5, the query was considered well-formed. This process ensures the query effectively captures diverse, relevant information and can help teach smaller models how to retrieve supporting evidence for answering questions.
@@
Average Answer Length (Word) 1.95
Average Source Documents Per Question 2.29
@@
4 Proposed Benchmark: M2QA-Bench
We introduce M2QA-Bench, a benchmark of 1k diverse Q&A pairs to evaluate LVLMs on complex MMQA with full documents. Unlike templated datasets (Talmor et al., 2021), questions are varied and challenging (see Appendix G for details on complexity and diversity). Answering requires cross-modal reasoning and information extraction from full documents, including images and tables. See Table 2 for key statistics (more in Appendix G) and Appendix I for samples generated by FM2DS for M2QA-Bench. To create this benchmark, we used the FM2DS pipeline to generate 1,200 samples, which were verified by three annotators for correctness, multihop reasoning, multimodality, and answer accuracy. Each sample was scored 1 (valid) or 0 (invalid). This annotation required minimal human effort (2.2 min/question on average) due to structured queries. Samples averaging below 0.75 were removed, leaving 1,142 (i.e removing only 5% of the total); we then randomly selected 1,000 for the benchmark to ensure consistency in evaluation and reduce potential sampling bias. Annotator agreement (Fleiss’ Kappa (Fleiss, 1971)) was 0.83.